Find the enthalpy of Helium if its internal energy is 200 KJ/kg


h = _____ kJ/kg


1
Expert's answer
2021-09-12T19:21:42-0400

Gives

U=200KJ/kg

"U=mc_vT\\H=mc_pT"

Now

"\frac{H}{U}=\frac{mc_pT}{mc_vT}\\\frac{H}{U}=\frac{c_p}{c_v}"

Helium is monoatomic

"\frac{c_p}{c_v}=\frac{5}{3}"


"H=\frac{c_v}{c_p}\times U=\frac{5}{3}\times200=333.33kJ/kg"

"h=H=333.33kJ\/kg"